OpenStudy (anonymous):

What is the area of a square window that has a perimeter of 20 feet?!!!!

jimthompson5910 (jim_thompson5910):

the perimeter is 20, so what is the side length?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

let x=side length per=4x 4x=20 x=5

jimthompson5910 (jim_thompson5910):

good, 20/4 = 5

OpenStudy (jdoe0001):

so you have a perimeter of 20, a flat unit so to make it square unit, borrow a few units from the 20, or split it in 2 use one unit for the length, another for the width and get the area by length x width

jimthompson5910 (jim_thompson5910):

the area is then found by using this formula A = s^2 A = area s = side length

OpenStudy (anonymous):

5 isn't right @jim_thompson5910 @ranjeet619

jimthompson5910 (jim_thompson5910):

5 is a step to get the answer, it's not the actual answer

OpenStudy (anonymous):

.... what do you do next then

OpenStudy (anonymous):

area is lengthxwidth, since its a square that has 4 equal sides do 20/4 to get 5. so each side is 5. lengthxwidth is 5x5 = 25

OpenStudy (theeric):

@xoxox123 , jim_thomspson5910 's response and warningiflirt 's response will help you see how to get the answer.

OpenStudy (anonymous):

I tried 25 already and it was wrongg

OpenStudy (theeric):

25 feet squared should be the right answer! Make sure the units are square feet.

OpenStudy (theeric):

Does it require the units?